1.      A study of 40 SHS English teachers established that they spent 12.6 minutes revising a student’s term paper.

a. Find the 90% confidence interval in the mean time for all composition papers when minutes.

b. If the teacher affirmed that she spent an average of 30 minutes revising a term paper, what would be your reaction?


1
Expert's answer
2022-04-05T06:49:08-0400

"CI=\mu \plusmn z \frac{\sigma}{\sqrt{n}}"

Z(90%)=1.64

n=40

"\mu=12.6"

"CI=12.6 \plusmn 1.64 \frac{\sigma}{\sqrt{40}}=12.6 \plusmn 0.26 \sigma"

b. "30=12.6+0.26 \sigma"

"\sigma=(30-12.6)/0.26=66.9"

If "\sigma" smaller then 66.9 min, then 30 min is too long.

If "\sigma" larger then 66.9, then it is ok.